OpenStudy (anonymous):

Factor Completely: 648x^3+375 Please show me a step by step way to do this problem so I can compare my work.

OpenStudy (anonymous):

@OrionsBelt

Directrix (directrix):

648x³ +375 = Factor out common monomial of 3 3(216x³ + 125)= Notice that 216x³ and 125 are perfect cubes 3 * ( [6x]³ + 5³ ) = Then, use the attached factoring pattern for the sum of two cubes 3 (6x + 5) ( 36x² - ? + ? ) @Mhm120